OpenStudy (anonymous):

What is the solution of log_2x + 6 8 = 3? @inkyvoyd

OpenStudy (anonymous):

@rm152

OpenStudy (anonymous):

Is it \[\log_2{x} +68=3\]?

OpenStudy (inkyvoyd):

assuming what @rm152 said, you want to get the log expression on one side of the equation...

OpenStudy (anonymous):

no it just says that all together rm152

OpenStudy (inkyvoyd):

so @woohoo is what @rm152 what you are trying to solve?

OpenStudy (anonymous):

x = −2 x = 2 x = −3 x = 3 these are the options

OpenStudy (inkyvoyd):

woohoo, can you try retyping the equation?

OpenStudy (inkyvoyd):

or maybe taking a screenshot of it...

OpenStudy (confusionist):

Is there any kind of sign between the 6 & 8?

OpenStudy (anonymous):

What is the solution of \[\log_{2x+6}8=3 \]

OpenStudy (anonymous):

no there isnt

OpenStudy (inkyvoyd):

OOH.

OpenStudy (confusionist):

8=(2x+6)^3 2x+6=2 x=2

OpenStudy (inkyvoyd):

skew it, I forgot why I quit helping people in the first place hehe

OpenStudy (anonymous):

Inkyvoyd you are doing good, if you can't help at least you can learn easier ways to solve

OpenStudy (confusionist):

Once we got a question we were on a roll, though!

OpenStudy (inkyvoyd):

@rm152 it has more to do with my slow tex typing speeds.

OpenStudy (confusionist):

I'm slowly improving on mine myself. x_x You beat me to getting him the answer, though! >:o

OpenStudy (anonymous):

Haha I can't even upload a picture in the right orientation.Typing takes too long!

OpenStudy (anonymous):

lol i thought i broke my neck when you posted that last picture

OpenStudy (ranga):

2x + 6 = 2 So x = -2 (not +2)

OpenStudy (anonymous):

You are right @ranga I put 2 and got it wrong